OpenStudy (hellomynameiskam):

Solve the equation. 1/4 + w = 6 1/3 A. w = 6 1/12 B. w = 6 1/7 C. w = 6 2/7 D. w = 6 7/12 @KendrickLamar2014

OpenStudy (kendricklamar2014):

1. Simplify both sides: \[\frac{ 1 }{ 4 } + w = \frac{ (6)(1) }{ 3 }\] \[= w + \frac{ 1 }{ 4 } = 2\]

OpenStudy (kendricklamar2014):

2. Subtract 1/4 from both sides:\[w + \frac{ 1 }{ 4 } - \frac{ 1 }{ 4 } = 2 - \frac{ 1 }{ 4 }\] \[w = \frac{ 7 }{ 4 }\]

OpenStudy (kendricklamar2014):

D is the answer

OpenStudy (hellomynameiskam):

7 divided by 4 is 1.75 or is it 4 divided by 7

OpenStudy (hellomynameiskam):

ohh okay. sorry didnt see the answer u put. (: